Systems Staff Meeting Notes 7/17/00

Systems Staff Meeting Notes
July 17, 2000

Present: Will Chaney, Bill Covey, Debra Harris, Winston Harris, Suzy Shaw, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. This is Geoffrey's first staff meeting!
  2. Ying's replacement has been named: Cynthia Bowker. She should arrive in Systems sometime within the next two weeks, although she will not be full time until September.
  3. Chris Leopold will be a guest at the Systems Staff Meeting on July 31. Chris is from NERDC and will be the backup reserve for Bill during his leave of absence.
Review Minutes of 7/3/00
  1. RIR #1 - The Preservation server has arrived. We'll need to figure out how to proceed.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. We'll be buying a few more licenses for NT.
  2. Geoff has a couple of replacement monitors due.
Liaisons
  1. Bill sent out a notice about the patch for O2K. SR1A has to be in place prior to download. It's in Read Only on Smathers NT2.
  2. IE 5.5 update is also in Read Only. Should be a fast update.
  3. FrontPage - people were having a problem saving to the server in FP2K. It was a known MS bug, caused by the SR1A patch. If users are experiencing this trouble, they need to put in 1.1 after the SR1A patch.
Networking
  1. Will came in yesterday and switched the East tower network feed. Everything went smoothly. This whole building should have a 100M feed.
  2. Last week, Will switched the West fiber feed to the Com Closet in room 229.
  3. Winston moved the West NT servers in Library Office to the Com Closet in room 219.
  4. Strangeness on the DLC move - Brad from Facilities showed up in Systems asking about removing network wiring on the second floor. Systems had received no notification of these plans. They wanted to remove poles and pods from the area. Network wiring won't work that way. We have to reach some sort of agreement with them.
  5. The Smathers CD ROM server is waiting for Tom Caswell to make the shortcuts work.
  6. 250 sign-on limit on Smathers PS - it keeps filling up and needs to be cleared out at least once a day. We're trying to get Xerox off of Novell, but that will probably take awhile - probably not by the fall term. We could possibly see about automating the clearing out process. Geoffrey will try to get the addresses and track down where the 'not logged on' entries are coming from.
  7. Need to finalize the DLC terabyte server configuration.
  8. Closet micros for troubleshooting - Geoffrey is working on it.
  9. Documentation of current net connections - Suzy and Bill will be working on Web-enabled diagrams of Library wiring at the switch level. It will provide a quick overview for NERDC. Would not be used to track down work stations. Will probably use VISIO software.
  10. The Digital crew is still working in Chemistry. Will is awaiting an estimate for the work to be done in Documents. He could not find Bobbie's notes on the wiring in West - will wait for her to return next week.
Software Projects Status
  1. Preservation Q&A - waiting for evaluation of data.
  2. Registry migration - complete; John Freund has the master version.
  3. Personnel system mods/database - revamping OPS upcoming, all current complete.
  4. OPS tracking - in progress.
  5. Library Statistics - first phase is complete. Phase two is in progress.
  6. Hilton is looking at VISIO for the Systems wiring database.
  7. Betsy Simpson is having trouble with her Chinese-Japanese-Korean software and the FCLA server. The problems were researched we've determined what we're doing wrong. Making progress towards resolution.
  8. Mellon 2 - no word on status in a long time, we'll need to find out where it went.
  9. DLC Web interfaces - Ying is working on a tracking database for Erich. He wants to track all of the DLC projects by cost, staff, hours, etc. So far they've had two meetings, trying to define the structure. It is all at the specs stage. Ying plans to use Delphi with the SQL server and a Web application for reports.
Calendar
  1. Bill will be out July 18 & July 21 afternoon.
  2. Winston out July 21.
  3. Bobbie is on vacation until July 24.
  4. Rich returns from vacation on July 26.
  5. Will out July 24-28.


Back to the Systems Homepage

Last updated July 24, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu